A former has to go 6m due north 8m due east and 10m upwards to resch his feild
what is distance  ?
and displecement?

Grade:12th pass

2 Answers

Manish Sah
3 years ago
These three displacements are mutually perpendicular to each other so                                Net disp.(s)=(8^2+6^2+10^2)^1/2=10(2)^1/2m
 
Distance=6+8+10=24m
Vikas TU
3 years ago
First, he goes 6m north, then 8m east and finally 10m north again. So we draw an imaginary triangle to find the displacement. Displacement is the shortest distance between the final point and the initial point. So the shortest distance= √(16^2 + 8^2) = √320= 17.88m or 18m ( approximately)